The world's longest suspension bridge is the akashi kaikyo bridge in Kobe Japan. It is constructed of steel (12*10-6°c) and has a total length of 3910 m. Find the change on length of the bridge from a cold day in winter (2.00°c) to a hot day in summer (35.0°c)


1
Expert's answer
2021-02-09T03:46:36-0500

Initial length of bridge L0= 3910m

Temp on cold day = 2⁰C

Temp on hot day = 35⁰C

Change in Temp ΔT = 35 - 2 = 33⁰C

"\alpha" = 12×10-6 ⁰C

Now find Change in length

ΔL = L0 "\alpha" "\Delta"T

= 3910 × 12 × 10-6 × 33

"\Delta L" = 1.54836m